OpenStudy (anonymous):

determine whether the equation defines y as a function of x? x^2 + 8y^2=1

OpenStudy (anonymous):

can someone help me?

OpenStudy (anonymous):

thanks!

OpenStudy (turingtest):

The equation as it is is not defined in terms of x explicitly. Are we allowed to change it?

OpenStudy (turingtest):

...as you have the equation right now it isn't in the form y=(something) so are we allowed to change it to answer your question? I'm unclear.

OpenStudy (earthcitizen):

\[ x^2 + 8y^2=1\]

OpenStudy (earthcitizen):

solve implicitly

OpenStudy (earthcitizen):

\[(dy(x))/(dx) = -x/(8 y)\]

OpenStudy (earthcitizen):

\[(dx(y))/(dy) = -(8 y)/x\]

OpenStudy (turingtest):

there is no derivative in this problem @earth

OpenStudy (earthcitizen):

it looked like an implicit function

OpenStudy (turingtest):

if y is supposed to be defined in terms of x it isn't. if we have to solve for y we wind up taking a +/- square root, so th graph would fail the vertical line test. it looks to me like that is all needed to say that this is not a function. at least not as it is.

OpenStudy (earthcitizen):

is ut a circle equation ?

OpenStudy (turingtest):

right, and a circle fails the vertical line test so no, not a function I say.

OpenStudy (turingtest):

sorry, not a circle, an ellipse

OpenStudy (turingtest):

still not a function

jhonyy9 (jhonyy9):

- so i think it like : y=f(x)=+/- sqrt((1-x2)/8)